A charge, q1, of +16 micro-coulombs is located at x = 0, y = 6.4 cm, a charge q2 of -40 micro-coulombs is located at x = 0, y = -2.8 cm, an unknown charge of q3, is located at x =- 3.3 cm, y = 0 cm, and a charge, q4, -6 micro-coulombs is located at x = +16.6 cm, y = 0 cm. The direction of the total electric force on q4 is 136.5 degrees from the +x axis, counter-clockwise. What is the unknown charge, q3, in micro-coulombs? If negative, include a negative sign in your answer.

Please explain what you are doing!
Answer #1

16 x 10 C 0.064 m F2 0.033 m 0.166 m 0 Q3-? .028m 4-6 x 104 c 89 40 x 10c °С

F = Net force on charge q4

Fx = x-component of the net force = F Cos136.5 = - 0.725 F

Fy = y-component of the net force = F Sin136.5 = 0.688 F

heta _{1} = tan-1(0.064/0.166) = 21.1 deg

heta _{2} = tan-1(0.028/0.166) = 9.6 deg

r1 = distance of charge q1 from q4 = sqrt(OA2 + OD2) = sqrt((0.064)2 + (0.166)2) = 0.178 m

r2 = distance of charge q2 from q4 = sqrt(OB2 + OD2) = sqrt((0.028)2 + (0.166)2) = 0.168 m

r3 = distance of charge q3 from q4 = 0.033 + 0.166 = 0.199 m

F1 = force by charge q1 on q4 = k q1 q4/r12 = (9 x 109) (16 x 10-6) (6 x 10-6)/(0.178)2 = 27.3 N

F2 = force by charge q2 on q4 = k q2 q4/r12 = (9 x 109) (40 x 10-6) (6 x 10-6)/(0.168)2 = 76.5 N

F3 = force by charge q3 on q4 = k q3 q4/r32 = (9 x 109) q3 (6 x 10-6)/(0.199)2 = (1.4 x 106) q3

Net force along the Y-direction is given as

Fy = F1 Sinheta _{1} + F2 Sinheta _{2}

Fy = (27.3) Sin21.1 + (76.5) Sin9.6

Fy = 22.6

0.688 F = 22.6

F = 32.85 N

force along the x-direction is given as

Fx = - F1 Cosheta _{1} + F2 Cosheta _{2} - F3

- 0.725 F = - F1 Cosheta _{1} + F2 Cosheta _{2} + F3

- (0.725) (32.85) = - (27.3) Cos21.1 + (76.5) Cos9.6 - (1.4 x 106) q3

q3 = 52.7 x 10-6 C